OpenStudy (anonymous):

hey, can somebody please help me? i have a HUGE text tomorrow! how can i find the roots of 3x^4+11x^3+14x^2+7x+1=0? i think this particular equation requires the quadratic formula also along the way all help is very much appreciated!! thank you!!!

OpenStudy (anonymous):

First thing to do is to get the possible roots of the equation. To wit \[3x^4+11x^3+14x^2+7x+1=0\]Get the factors of the coefficient of the term of highest degree (here, \(x^4\)). The factors of 3 are \(\pm 1 \text{ and } \pm3\). Now, get the factors of the constant term, (here, 1). The factors of 1 are \(\pm{1}\). To put it altogether: \(p\): (for constant term) \(\pm{1}\) \(q\): (for the \(x^4\) term) \(\pm{1}, \pm{3}\) Therefore: \(\frac{p}{q}\): (possible roots of the equation) \(\pm{1}, \pm{\frac{1}{3}}\). Now use synthetic division to test each possible root.

OpenStudy (anonymous):

thank you so much @Jnlucero!! but i thought we had to use the quadratic formula in this one?

OpenStudy (anonymous):

Nope, you have to do synthetic division first to find some of the roots. When the reduced equation becomes quadratic, that's the time we have to use either factoring or quadratic formula.

OpenStudy (anonymous):

can you do me a huge favor and do this one so i can see an example?

OpenStudy (anonymous):

@Jnlucero

OpenStudy (anonymous):

sure. can you wait for a while?

OpenStudy (anonymous):

yes, absolutely, thank you so much!!

OpenStudy (anonymous):

_-1_| 3 11 14 7 1 try -1 __________-3___-8___-6__-1 3 8 6 1 0 Since the last number is zero, then -1 is a root. Now consider only the first four numbers. _-1_| 3 8 6 1 try -1 again (yes, it can repeat many times) __________-3___-5___-1 3 5 1 0 Again, the last number zeroes, so -1 is again a root. Now consider only the first three numbers. This means the reduced equation is \(3x^2+5x+1=0\), which can be solved using quadratic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-5 \pm \sqrt{5^2-4(3)(1)}}{2(3)}=\frac{-5\pm\sqrt{13}}{6}\]So, we have obtained the two other roots.

OpenStudy (anonymous):

so what would the final answer be?

OpenStudy (anonymous):

and thank you sooooo so so much, you've been super helpful! xx

OpenStudy (anonymous):

Just collect the results. x = -1 (double root), x = \(\frac{-5\pm\sqrt{13}}{6}\). Your very much welcome. :)

OpenStudy (anonymous):

so it would be -1 multiplity 2, and then the fraction

OpenStudy (anonymous):

yep. And one thing, its "multiplicity".